### v1

# Theoretical Foundation of Quantum Stochastic Contrastive Learning for WiFi-Based Indoor Localization with Floor and Building Detection

Abstract
WiFi-based indoor localization is essential for applications like asset tracking, healthcare monitoring, and smart buildings. However, existing systems face challenges such as data variability, environmental noise, and difficulty detecting floor and building levels, compounded by limited labeled data and high costs of RSS collection. This paper introduces quantum stochastic contrast learning (QSCL), a novel framework grounded in rigorous theoretical foundations. We present four theorems and one lemma that establish the probabilistic augmentation, diversity enhancement, relationship preservation, and resilience of QSCL under quantum noise, supported by formal proofs. Leveraging these foundations, QSCL utilizes quantum computing to generate strong data augmentations with stochastic perturbations, enhancing data diversity, while classical weak augmentations provide subtle variations for robust feature learning. To capture spatial and temporal dependencies in sequential data, we propose a spatial temporal adaptive attention (STAA) encoder that integrates convolutional layers with adaptive attention mechanisms. Furthermore, a bidirectional contrastive loss function is introduced to capture forward and reverse relationships between augmented views, ensuring robust representations. Comprehensive evaluations on the UJIIndoorLoc and UTSIndoorLoc datasets validate QSCL, demonstrating superior performance with reduced labeled data and resilience to quantum noise such as bit-flip, dephasing, and measurement noise. The proposed framework significantly improves localization accuracy, floor and building detection, and generalizability in challenging indoor environments.
Index Terms:
WiFi indoor localization, quantum stochastic learning, contrastive learning representation, quantum augmentation, flood and building detection, quantum computing, received signal strength (RSS).
I Introduction
Indoor localization has emerged as a critical area of research and development due to the exponential growth of mobile devices and the increasing demand for location-based services (LBS). An accurate indoor location is essential for applications such as healthcare care, asset tracking, smart factories, and emergency response [Sridharan2020, Wyffels2014, Zhang2014, Wang2024]. Unlike outdoor systems that rely on GPS, indoor environments pose unique challenges, including signal attenuation and multipath interference, which render GPS ineffective [Luo2019].
To address these challenges, researchers have increasingly adopted WiFi-based fingerprinting as a cost-effective solution that leverages existing infrastructure [Xu2012]. This approach constructs unique location fingerprints using the received signal strength indicator (RSSI) of access points (AP), allowing position estimation during the online phase [Kumar2016]. However, RSS measurements are highly sensitive to environmental factors, such as human movement, structural changes, and signal interference, leading to significant variability that undermines the precision of localization [He2016, Li2021]. Thus, addressing noise and variability is critical for developing reliable WiFi-based indoor localization systems [Chen2023].
Machine learning and deep learning methods, including convolutional neural networks (CNNs) and recurrent neural networks (RNNs), have been employed to model RSS patterns and enhance localization performance [Zheng2021, He2023]. More recently, graph neural networks (GNNs) have been utilized to capture spatial relationships inherent in RSS data, offering improved robustness in dynamic and noisy environments [Chen2023]. Despite these advancements, existing methods still face challenges in generalizability and handling data variability, particularly under noisy and unseen conditions [He2023].
Supervised learning approaches are often limited by their reliance on labeled datasets, which are scarce in real-world scenarios. Contrastive learning, using unlabeled data to learn robust feature representations, has emerged as a promising alternative. By identifying similarities and differences within the data, contrastive learning improves generalization under limited supervision [Chen2020, Khosla2020]. However, current contrastive learning techniques struggle with effective augmentation strategies and fail to model the complex spatio-temporal dependencies needed to address the stochastic nature of noisy environments.
In parallel, the emergence of quantum computing offers transformative potential to address these challenges. Quantum properties, such as superposition and entanglement, enable efficient computations that surpass the capabilities of classical approaches [Preskill2018, Arute2019]. In machine learning, quantum algorithms can accelerate training, optimize feature extraction, and model complex data relationships more effectively [Biamonte2017]. Integrating quantum-assisted learning into classical frameworks not only enhances generalization, but also provides a robust foundation for effective augmentation and representation learning in noisy environments [Gao2022, Cao2019].
To address the challenges of data variability, environmental noise, and limited labeled data in WiFi-based indoor localization, advances in augmentation strategies and spatiotemporal encoding are essential. Stochastic augmentations introduce the diversity required for robust learning under variable conditions, while enhanced spatio-temporal models effectively capture dynamic relationships within sequential data. In addition, a well-designed contrastive loss function ensures meaningful and generalizable feature representations.
This paper introduces quantum stochastic contrast learning (QSCL), a novel framework that combines rigorous theoretical foundations with empirical validation to advance WiFi-based indoor localization. QSCL is grounded in four theorems and one lemma, supported by formal proofs, which establish the probabilistic diversity of quantum augmentations, resilience to quantum noise, and preservation of meaningful relationships in augmented data. The framework employs quantum-driven strong augmentations alongside classical weak augmentations to enhance robustness and generalization. A Spatio-Temporal Adaptive Attention (STAA) encoder is utilized to capture spatial and temporal dependencies, while a bidirectional contrastive loss ensures effective representation learning. Empirical evaluations on benchmark datasets, including UJIIndoorLoc [UJIIndoorLoc2014] and UTSIndoorLoc [UTSIndoorLoc2023], validate QSCL’s superior performance in localization accuracy, distance error estimation, and robustness to noise. By addressing real-world challenges such as data variability and limited labeled data, QSCL provides a scalable and effective solution for practical WiFi-based indoor localization applications, including floor and building detection.
The structure of the paper is as follows. Section II reviews related work. Section III presents a detailed methodology of the framework. In Section IV, we present the impact of quantum noise. Section V provides experimental results and discussions. Finally, Section VI concludes the paper.
II Related Work
We evaluate the existing research on indoor localization, organizing it into four main categories: 1) techniques for fingerprint-based indoor localization, 2) approaches to reduce RSS uncertainty, 3) strategies for AP deployment, and 4) Baseline.
II-A Fingerprint-Based Indoor Localization Techniques
For many years, indoor localization research has heavily relied on fingerprint-based techniques. To enhance the accuracy of received signal strength (RSS) fingerprinting, conventional machine learning algorithms like sequential minimum optimization [Varshavsky2007], decision trees [Bozkurt2015], and k-nearest neighbors (KNN) [Apostolo2019] have been extensively utilized. In the past few years, deep learning has emerged as a promising approach for advancing indoor positioning, owing to its excellent ability for feature representation. Kim et al. [Kim2018] employed stacked autoencoders and feed-forward deep neural networks (DNNs) for floor classification tasks, achieving notable outcomes in extensive indoor areas. Song et al. [Song2019] developed CNNLoc, a one-dimensional convolutional neural network that enhanced both floor classification and regression performance. Similarly, Cha and Lim [Cha2022] introduced a hierarchical auxiliary deep neural network (HADNN), which consists of consecutive feed-forward neural networks (CFNNs), to efficiently manage complex indoor conditions. Furthermore, recurrent neural networks (RNNs) have been adopted to capture temporal features for scenarios with dynamic user movement. Hoang et al. [Hoang2019] applied long short-term memory (LSTM) and bidirectional LSTM models to improve trajectory tracking and localization precision.
II-B Solutions to Mitigate RSS Uncertainty
RSS-based localization is prone to substantial uncertainty owing to environmental influences such as multipath propagation and signal attenuation. Initial approaches, like the signal propagation models applied by Xu et al. [Xu2012], sought to forecast the spatial distribution of signals for radio map updates. Nevertheless, these models frequently encountered difficulties in handling intricate indoor settings. To overcome these challenges, He et al. [He2016] introduced a crowdsourcing method where devices functioned simultaneously as “surveyors” and “queriers” to constantly refresh the radio map. Kumar et al. [Kumar2016] applied Gaussian process regression to fill in RSS values, thereby decreasing the manual work necessary for keeping the radio map current. More recently, graph neural networks (GNNs) have been leveraged for their effectiveness in incorporating spatial relationships. Zheng et al. [Zheng2021] utilized heterogeneous GNNs to grasp the spatial closeness between access points (APs) and reference points (RPs), thus mitigating RSS fluctuation. In addition, He et al. [He2023] used geomagnetic signals as graph nodes and RSS as connections to form adaptive graph models, using graph attention networks for joint indoor localization.
II-C AP Deployment Strategies
The strategic placement of access points (APs) is vital for enhancing localization precision. Conventional approaches typically aim to extend signal coverage [Du2017]. For instance, Liao et al. [Liao2011] utilized signal coverage thresholds to fine-tune AP placement, whereas Du and Yang [Du2017] implemented particle swarm optimization algorithms to optimize the alignment of fingerprints at reference points, facilitating efficient AP positioning. Luo et al. [Luo2019] investigated the variability of RSS data to determine the best access points (APs). Yet, methods that prioritize signal reach and fingerprint likeness do not always result in precise localization [Du2017, Zirazi2012]. Zirazi et al. [Zirazi2012] tackled this problem by reducing the geometric dilution of precision, thus improving the accuracy of newly installed APs. Tian et al. [Tian2020] suggested a comprehensive search strategy to find the most effective AP deployment solutions; nevertheless, this method encounters scalability challenges as the number of APs rises.
II-D Baseline
Wang et al. [BaselineGNNWiFi_Wang2024] present an end-to-end neural network for AP selection, designed to optimize access point usage while ensuring high localization precision, thus offering a scalable approach for dynamic indoor settings. Despite its promise, the method faces significant challenges, such as elevated hardware and storage costs due to the large number of APs required, and sensitivity to RSS fluctuations from environmental changes. The dependency on RSS for localization presents obstacles in scalability and adaptability to fast-changing environments, affecting real-time accuracy and durability. Moreover, the model’s performance may suffer when ground truth data is sparse or nonexistent, limiting its reliability in real-world applications.
Figure 1: Quantum Stochastic Contrastive Learning for WiFi-Based Indoor Localization
III Method
In this section, we present the methodology behind the proposed quantum-assisted augmentation with contrastive learning framework for WiFi-based indoor localization. Our approach combines quantum computing techniques with classical machine learning methods to address the challenges posed by data variability, environmental noise, and limited labeled datasets. In the following section, we discuss this in detail.
III-A Problem Formulation and Objective
WiFi-based indoor location is based on the received signal strength indicator (RSSI) values of multiple access points (AP) to estimate the position of a device and classify its floor and building within a given structure [Zafari2019, Liu2020WiFi]. Given unlabeled data for the WiFi signal $X=\{x_{1},x_{2},\dots,x_{N}\}$, where each $x_{i}\in\mathbb{R}^{d}$ represents a feature vector of RSSI values from the access points $d$, the objective is to learn a mapping ($f:X\to Y$) which predicts $Y=(y_{\text{floor}},y_{\text{building}},y_{\text{position}})$ as shown in Eq. 1.
$Y=(y_{\text{floor}},y_{\text{building}},y_{\text{position}})\in\mathbb{R}^{3}.$ (1)
The proposed framework employs quantum-assisted augmentation to introduce diversity via strong augmentations using quantum noise operators $\mathcal{N}_{\text{quantum}}(x_{i})$, producing quantum-augmented views $\mathcal{A}_{s}(x_{i})$ as shown in Eq. 2.
$\mathcal{A}_{s}(x_{i})=\mathcal{N}_{\text{quantum}}(x_{i}).$ (2)
To enhance learning, both quantum-assisted strong augmentations $\mathcal{A}_{s}$ and classical weak augmentations $\mathcal{A}_{w}$ are applied, generating views $\tilde{x}_{i}$ and $\hat{x}_{i}$ as shown in Eq. 3.
$\tilde{x}_{i}=\mathcal{A}_{s}(x_{i}),\quad\hat{x}_{i}=\mathcal{A}_{w}(x_{i}).$ (3)
The learning objective maximizes the consistency of representations in augmented views using a similarity function $\text{sim}(\cdot,\cdot)$ and an encoder $f_{\theta}(\cdot)$, formalized as shown in Eq. 4.
$\max_{\theta}\mathbb{E}\left[\text{sim}(f_{\theta}(\tilde{x}_{i}),f_{\theta}(\hat{x}_{i}))\right].$ (4)
The contrastive loss function, shown in Eq. 5, is designed to minimize intra-cluster similarity while maximizing inter-cluster separation.
$\min_{\theta}\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(f_{\theta}(x_{i}),f_{\theta}(x_{j})))}{\sum_{k=1}^{N}\exp(\text{sim}(f_{\theta}(x_{i}),f_{\theta}(x_{k})))}.$ (5)
Intra-cluster and inter-cluster distances are defined in Eq. 6 and Eq. 7, respectively.
$d_{\text{intra}}(x_{i},x_{j})=\|f(x_{i})-f(x_{j})\|^{2},$ (6)
$d_{\text{inter}}(x_{i},x_{k})=\|f(x_{i})-f(x_{k})\|^{2}.$ (7)
The objective minimizes the ratio of intra-cluster to inter-cluster distance, ensuring robust feature separation, as shown in Eq. 8.
$\text{Minimize}\quad\frac{d_{\text{intra}}}{d_{\text{inter}}}.$ (8)
This approach effectively enhances localization accuracy, noise resistance, and generalization in multi-floor and multi-building indoor environments.
Figure 2: Quantum-Based Augmentation Circuits
III-B Strong Quantum Stochasticity Augmentation
The received signal strength indicator (RSSI) $X$, widely used for indoor localization [Zafari2019, Liu2020WiFi], is scaled in rotation angles to enable robust and generalizable representations [Biamonte2017, Schuld2015, Chen2020]. The rotation angle $\theta_{i}$ for the $i$-th RSSI value is calculated as shown in Eq. 28.
$\theta_{i}=\frac{X_{i}\cdot\pi}{\max(X)},$ (9)
where $X_{i}$ is the RSSI value for the $i$-th access point (AP), and $\max(X)$ is the maximum RSSI value in the dataset. These angles parameterize quantum rotation gates $R_{x}(\theta)$, defined as shown in Eq. 15.
$R_{x}(\theta)=e^{-i\frac{\theta}{2}\sigma_{x}}=\begin{pmatrix}\cos\frac{\theta}{2}&-i\sin\frac{\theta}{2}\\ -i\sin\frac{\theta}{2}&\cos\frac{\theta}{2}\end{pmatrix},$ (10)
where $\sigma_{x}$ is the Pauli-X operator and shown in algorithm 1. This process constitutes the strong augmentation, which is performed using a quantum circuit $R_{\alpha}$, parameterized by $\alpha$. The circuit applies quantum operations such as rotations and entanglement to transform the input RSSI signal $X$ into an enriched quantum representation. The augmented version $h_{1}^{S}$ is given by:
$h_{1}^{S}=R_{\alpha}\left(\frac{X_{i}}{\max(X)}\right)q_{i},$ (11)
where $q_{i}$ represents the resulting quantum state after applying the circuit. The normalization term $\frac{X_{i}}{\max(X)}$ ensures that the input data is scaled appropriately for quantum processing, maintaining numerical stability and consistency.
Input: RSSI values $X=\{X_{1},X_{2},\dots,X_{n}\}$ from $n$ access points
Output: Quantum-augmented view $X_{\text{aug}}$
Step 1: Normalize RSSI values
for
 $X_{i}\in X$ do
       Compute rotation angle: $\theta_{i}=\frac{X_{i}\cdot\pi}{\max(X)}$
end for
Step 2: Apply Quantum Circuit
for
 $\theta_{i}$ do
[…]
end for
Step 3: Measure Quantum State
for
[…]
       Measure the qubit and store the result: $X_{\text{aug}}\leftarrow$ measurement outcome
end for
return $X_{\text{aug}}$
Algorithm 1 Quantum-Based Augmentation of RSSI Values
III-B1 Theoretical Foundation with Theorems and Proof
The proposed framework leverages quantum-based augmentation to address challenges in data variability, noise, and generalization for WiFi-based indoor localization. Grounded in key theoretical principles, the framework ensures robustness through probabilistic augmentations (Theorem III.1), diversity via quantum noise and measurement randomness (Theorem III.2), and semantic consistency by preserving scaled relationships (Theorem III.3). We will discuss in detail as follows:
Theorem III.1.
(Probabilistic Augmentation): Quantum rotation gates probabilistically augment RSSI values with bounded stochastic deviations from the original values.
$X^{\prime}_{i}=X_{i}+\delta,\quad\text{where }|\delta|\leq\frac{\pi}{2}.$ (12)
Proof.
Let $\theta_{i}$ be the rotation angle corresponding to a normalized RSSI value $X_{i}$, where:
$\theta_{i}=\frac{X_{i}\cdot\pi}{\max(X)}.$ (13)
Each RSSI value $X_{i}$ is normalized to $\theta_{i}$ as in Eq. (13) and mapped to the quantum state:
[…]
The rotation gate $R_{x}(\theta_{i})$ is given by:
$R_{x}(\theta_{i})=e^{-i\frac{\theta_{i}}{2}\sigma_{x}}=\begin{pmatrix}\cos\frac{\theta_{i}}{2}&-i\sin\frac{\theta_{i}}{2}\\ -i\sin\frac{\theta_{i}}{2}&\cos\frac{\theta_{i}}{2}\end{pmatrix}.$ (15)
[…]
Since $\cos^{2}x+\sin^{2}x=1$, these probabilities are normalized.
During measurement, the qubit collapses to either:
[…]
The measurement outcome determines the augmented angle $\theta^{\prime}_{i}$:
[…]
Define the deviation $\delta$ as:
$\delta=\theta^{\prime}_{i}-\theta_{i}.$ (19)
From the measurement probabilities:
[…]
    $\delta\approx 0\quad(\theta^{\prime}_{i}\approx\theta_{i}).$ (20)
[…]
    $\delta\approx\pm\frac{\pi}{2}\quad(\theta^{\prime}_{i}\neq\theta_{i}).$ (21)
Thus, the perturbation $\delta$ is bounded:
$|\delta|\leq\frac{\pi}{2}.$ (22)
Reconstruct the augmented RSSI value $X^{\prime}_{i}$ from $\theta^{\prime}_{i}$:
$X^{\prime}_{i}=\frac{\theta^{\prime}_{i}\cdot\max(X)}{\pi}.$ (23)
Given $\theta^{\prime}_{i}\approx\theta_{i}\pm\delta$, we have:
$X^{\prime}_{i}\approx X_{i}\pm\frac{\delta\cdot\max(X)}{\pi}.$ (24)
The stochasticity introduced by quantum measurement ensures that the augmented RSSI value $X^{\prime}_{i}$ probabilistically deviates from the original value $X_{i}$, with the relationship:
$X^{\prime}_{i}\approx X_{i}\pm\frac{\delta\cdot\max(X)}{\pi}.$ (25)
This stochastic perturbation $\delta$ is critical for contrastive learning, as it generates diverse yet meaningful augmented representations in the feature space. ∎
Theorem III.2.
(Diversity Through Noise and Quantum Measurement): Quantum measurement randomness and noise ensure augmented values are diverse and distinct from the original data.
$X^{\prime}_{i}\neq X_{i},$ (26)
Proof.
The augmented value $X^{\prime}_{i}$ is given by:
$X^{\prime}_{i}=\frac{\theta^{\prime}_{i}\cdot\max(X)}{\pi}+\eta,$ (27)
where:
  • $\theta^{\prime}_{i}$ is derived from quantum measurement:
[…]
  • $\eta\sim\mathcal{N}(0,\sigma^{2})$ is a noise term.
The difference between the original RSSI value $X_{i}$ and the augmented value $X^{\prime}_{i}$ is:
$\Delta X_{i}=X^{\prime}_{i}-X_{i}=\frac{(\theta^{\prime}_{i}-\theta_{i})\cdot\max(X)}{\pi}+\eta.$ (29)
1. If $\theta^{\prime}_{i}=\theta_{i}$:
$\Delta X_{i}=\eta.$ (30)
Since $\eta\sim\mathcal{N}(0,\sigma^{2})$, the probability that $\Delta X_{i}=0$ is negligible, ensuring $X^{\prime}_{i}\neq X_{i}$ for most cases.
2. If $\theta^{\prime}_{i}\neq\theta_{i}$:
$\Delta X_{i}=\frac{(\theta^{\prime}_{i}-\theta_{i})\cdot\max(X)}{\pi}+\eta\neq 0.$ (31)
[…]
Thus, the combination of noise $\eta$ and quantum measurement randomness guarantees diversity, such that $X^{\prime}_{i}\neq X_{i}$ for most cases. ∎
Theorem III.3.
(Preservation of Relationships): Augmented RSSI values retain a meaningful scaled relationship with the original values, preserving semantic consistency.
$X^{\prime}_{i}=X_{i}+\Delta$ (32)
Proof.
The augmented RSSI value $X^{\prime}_{i}$ is defined as:
$X^{\prime}_{i}=\frac{\theta^{\prime}_{i}\cdot\max(X)}{\pi}+\eta,$ (33)
where $\theta^{\prime}_{i}$ is the augmented angle, and $\eta$ is a noise term. The relationship between $X^{\prime}_{i}$ and the original RSSI value $X_{i}$ can be expressed as:
$X^{\prime}_{i}=X_{i}+\Delta,$ (34)
where:
$\Delta=\frac{(\theta^{\prime}_{i}-\theta_{i})\cdot\max(X)}{\pi}+\eta.$ (35)
If $\eta=0$ and $\theta^{\prime}_{i}=\theta_{i}$, then:
$\Delta=0\implies X^{\prime}_{i}=X_{i}.$ (36)
Conversely, if $\eta\neq 0$ or $\theta^{\prime}_{i}\neq\theta_{i}$, then:
$\Delta\neq 0\implies X^{\prime}_{i}\neq X_{i}.$ (37)
The perturbation $\Delta$ is bounded by:
$|\Delta|\leq\frac{\max(X)}{\pi}\cdot|\theta^{\prime}_{i}-\theta_{i}|+|\eta|,$ (38)
ensuring that while $X^{\prime}_{i}$ deviates from $X_{i}$, the relationship remains meaningful and scaled. This establishes the bounded stochastic nature of the augmentation process, preserving the semantic relationship between the original and augmented values. ∎
These theorems, supported by rigorous proofs, are implemented through a hybrid approach combining quantum-driven strong augmentations with classical weak augmentations, enhancing robustness and generalizability in real-world scenarios.
III-C Weak Classical Augmentation
Weak augmentation is performed using a classical augmentation strategy that adds Gaussian noise to the original input data. This method results in $h_{2}^{W}$, defined and shown in eq 39.
$h_{2}^{W}=X+\mathcal{N}(0,\sigma^{2})$ (39)
where $\mathcal{N}(0,\sigma^{2})$ denotes Gaussian noise with zero mean and variance $\sigma^{2}$.
The combination of these two augmentation strategies allows the model to learn from both significant and minor perturbations, fostering a robust feature extraction process. The augmented WiFi signal data $h_{1}^{S}$ and $h_{2}^{W}$ are processed through separate base encoder, which extract meaningful representations.
III-D Encoder Networks
III-D1 Base Encoder
The base encoder architecture consists of a convolutional neural network (CNN) designed to extract robust feature representations from input sequences. The input data $X\in\mathbb{R}^{B\times I}$, where $B$ denotes the batch size and $I$ denotes the input dimension, is processed by a series of convolutional layers. The base encoder is defined and shown is eq 42.
[…]
where $W_{1}$ and $W_{2}$ are learnable parameters representing the weights of the convolutional filters, and AdaptiveMaxPool1D adjusts the output size to a fixed sequence length $L$. The resulting output $X_{\text{cnn}}\in\mathbb{R}^{B\times F\times L}$, with $F$ denoting the number of filters.
III-D2 Proposed SpatioTemporal Adaptive Attention Encoder
To capture both spatial and temporal dependencies inherent in sequential data, we propose a Spatio Temporal Adaptive Attention (STAA) encoder as shown in Fig 3. This encoder combines convolutional operations with adaptive attention mechanisms to learn complex representations effectively. The architecture is composed of a convolutional layer followed by a multi-head attention mechanism. The attention mechanism dynamically learns importance weights using adaptive scores. The convolutional layer processes the input sequence $X_{\text{cnn}}$ and outputs is shown in eq 43.
$X_{\text{conv}}=\text{Conv1D}(X_{\text{cnn}},W_{\text{conv}},\text{dilation\_rate}=d)$ (43)
where $W_{\text{conv}}$ represents the convolutional kernel weights, and $d$ denotes the dilation rate used for capturing long-range dependencies. Given the output $X_{\text{conv}}$, we apply a multi-head attention mechanism with adaptive weight scoring. The attention output $X_{\text{attn}}$ is computed as:
$X_{\text{attn}},\_=\text{MHA}(Q=X_{\text{conv}},K=X_{\text{conv}},V=X_{\text{conv}})$ (44)
where MHA is multihead attention and an adaptive scoring function $f_{\text{adapt}}$ computes attention weights as shown in eq 45.
$\alpha_{i}=\sigma(W_{\text{adapt}}\cdot\text{mean}(X_{\text{conv}}))$ (45)
where $W_{\text{adapt}}$ are learnable weights, $\sigma$ denotes the sigmoid activation, and $\alpha_{i}$ are the adaptive attention scores. The final output of the encoder is computed by combining convolutional and attention outputs:
$X_{\text{final}}=\text{LayerNorm}(X_{\text{conv}}+\alpha\otimes X_{\text{attn}})$ (46)
Figure 3: Proposed spatiotemporal adaptive attention encoder.
III-E Proposed Contrastive Loss Function
The SimCLR framework utilizes a contrastive loss function to maximize agreement between different augmentations of the same instance while minimizing agreement between augmentations of different instances [chen2020simple]. The loss is defined as:
$\mathcal{L}_{i,j}=-\log\frac{\exp(\text{sim}(z_{i},z_{j})/\tau)}{\sum_{k=1}^{2N}\mathbf{1}_{[k\neq i]}\exp(\text{sim}(z_{i},z_{k})/\tau)},$ (47)
where $\text{sim}(z_{i},z_{j})$ represents the similarity between latent representations $z_{i}$ and $z_{j}$, $\tau$ is the temperature parameter, and $N$ is the batch size.
The SimCLR loss focuses on unidirectional relationships, emphasizing positive pairs while contrasting with negatives, which limits its effectiveness in tasks with complex spatial-temporal dependencies like WiFi-based indoor localization. To overcome this, we propose a bidirectional contrastive loss that considers relationships in both directions ($i\to j$ and $j\to i$), enhancing feature extraction and representation learning. Encoder outputs are normalized via a projection head onto a unit hypersphere, ensuring consistency in similarity calculations. The final representations for strong and weak augmentations, $C_{st}$ and $C_{wt}$, are optimized with the proposed bidirectional contrastive loss:
$L_{STC}=\frac{1}{2}(L_{S}+L_{W}),$ (48)
where $L_{S}$ and $L_{W}$ are the losses for strong and weak augmentations:
$L_{S}=-\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(C_{st}^{i},C_{wt}^{i})/\tau)}{\sum_{j=1}^{B}\exp(\text{sim}(C_{st}^{i},C_{wt}^{j})/\tau)},$ (49)
$L_{W}=-\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(C_{wt}^{i},C_{st}^{i})/\tau)}{\sum_{j=1}^{B}\exp(\text{sim}(C_{wt}^{i},C_{st}^{j})/\tau)}.$ (50)
Here, $B$ is the batch size, $\text{sim}(\cdot,\cdot)$ denotes a similarity function (e.g., cosine similarity), and $\tau$ is the temperature parameter. The bidirectional loss ensures consistent and robust embeddings across augmented views, enhancing the model’s generalization and robustness.
Input: Unlabeled WiFi signal data $X=\{x_{1},x_{2},\dots,x_{N}\}$, labels $y=\{y_{1},y_{2},\dots,y_{N}\}$
Output: Fine-tuned model
Hyperparameters: Temperature $\tau$, Learning rate $\eta$, Batch size $B$, Epochs $E$
Initialize Base Encoder $\text{Enc}(\cdot)$, Spatio-Temporal Encoder $\text{STAA}(\cdot)$, and Projection Head
for epoch $e=1$ to $E$ do
       for mini-batch $\{x_{i},y_{i}\}_{i=1}^{B}$ do
             Step 1: Generate Augmented Data:
$h_{1}^{S}=R_{\alpha}\left(\frac{x_{i}}{\max(x)}\right)q_{i}$
(strong augmentation),
$h_{2}^{W}=x_{i}+\mathcal{N}(0,\sigma^{2})$ (weak augmentation). Step 2: Encode Representations
$Z_{st}=\text{Enc}(h_{1}^{S})$,  $Z_{wt}=\text{Enc}(h_{2}^{W})$
$C_{st}=\text{STAA}(Z_{st})$,  $C_{wt}=\text{STAA}(Z_{wt})$
Step 3: Compute Contrastive Loss
$L_{S}=-\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(C_{st}^{i},C_{wt}^{i})/\tau)}{\sum_{j=1}^{B}\exp(\text{sim}(C_{st}^{i},C_{wt}^{j})/\tau)}$
$L_{W}=-\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(C_{wt}^{i},C_{st}^{i})/\tau)}{\sum_{j=1}^{B}\exp(\text{sim}(C_{wt}^{i},C_{st}^{j})/\tau)}$
$L_{STC}=\frac{1}{2}(L_{S}+L_{W})$
Step 4: Update Weights
$\theta\leftarrow\theta-\eta\nabla_{\theta}L_{STC}$
       end for
end for
Fine-Tuning Phase:
Transfer learned weights to regression and classification heads
Regression Loss: $L_{\text{reg}}=\frac{1}{B}\sum_{i=1}^{B}(y_{i}-\hat{y}_{i})^{2}$
Classification Loss: $L_{\text{class}}=-\sum_{i=1}^{B}\sum_{c=1}^{C}y_{i}^{(c)}\log p_{\theta}(y_{i}^{(c)}|x_{i})$
Total Fine-Tuning Loss: $L_{\text{total}}=\lambda_{\text{reg}}L_{\text{reg}}+\lambda_{\text{class}}L_{\text{class}}$
Update Weights: $\theta\leftarrow\theta-\eta\nabla_{\theta}L_{\text{total}}$
Algorithm 2 Pre-Training Algorithm for Contrastive Learning
III-F Proposed Contrastive Learning Representation
The proposed methodology is outlined in Algorithm 2, with a visual depiction provided in Figure 1. This approach employs two augmentation $h_{1}^{S}$ (strong) and $h_{2}^{W}$ (weak) are processed through base encoder and STAA encoder to generate meaningful representations $C_{st}$ and $C_{wt}$. These embeddings are normalized onto a hypersphere, facilitating effective similarity calculations in the contrastive loss framework. The contrastive objective ensures robust embeddings across augmented views, improving model performance under noisy conditions. Following pre-training, the learned weights are fine-tuned for two tasks: distance regression and floor/building classification. The regression head minimizes mean squared error:
$L_{\text{regression}}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2},$ (51)
while the classification head minimizes cross-entropy loss:
$L_{\text{classification}}=-\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i}^{(c)}\log p_{\theta}(y_{i}^{(c)}|x_{i}).$ (52)
The total fine-tuning loss is defined as a weighted combination of regression and classification objectives:
$L_{\text{total}}=\lambda_{\text{reg}}L_{\text{regression}}+\lambda_{\text{class}}L_{\text{classification}},$ (53)
where $\lambda_{\text{reg}}$ and $\lambda_{\text{class}}$ are hyperparameters that balance the contributions of the regression and classification losses, respectively. This formulation enables the model to optimize both tasks effectively, ensuring balanced performance across regression and classification metrics. By integrating quantum-driven augmentations with Spatio-Temporal Adaptive Attention (STAA) Encoders, the framework achieves robust and generalizable representations, facilitating accurate and reliable indoor localization even under diverse and noisy conditions.
IV Impact of Quantum Noise on Quantum Stochasticity Augmentation
In this paper, we propose novel theoretical foundations to address the challenges posed by quantum noise in NISQ (Noisy Intermediate-Scale Quantum) devices, which inherently affect quantum operations. Our contributions include the introduction of Theorem IV.1 and Lemma IV.2, which lay the groundwork for understanding and mitigating the impact of quantum noise on quantum stochasticity augmentation.
Quantum noise, such as bit-flip, phase-flip, and depolarizing noise, affects the strong augmented view $C_{st}^{i}$, while the weak augmented view $C_{wt}^{i}$ remains unaffected. The combined effect of these noise types on the quantum state is represented as:
[…]
where noise operators modeled using Pauli matrices introduce stochastic perturbations. To ensure the robustness of the proposed augmentation framework, Theorem IV.1 proves that cosine similarity between the strong augmented view $C_{st}^{i}$ and the weak augmented view $C_{wt}^{i}$ remains consistent under quantum noise:
$\text{sim}(C_{st}^{i},C_{wt}^{i})=\text{sim}((C_{st}^{i}+\mathcal{N}_{st}),C_{wt}^{i}).$ (55)
This novel result guarantees that the augmented views retain their relationships, even when subject to noise, enabling reliable representation learning. Furthermore, Lemma IV.2 establishes that quantum noise operators, being stochastic and uncorrelated with the weak augmented view, introduce no systematic bias:
$\mathbb{E}[\mathcal{N}_{st}\cdot C_{wt}^{i}]=0.$ (56)
This lemma ensures that the expected contribution of noise to the similarity computation is null, preserving the integrity of the augmented views.
These theoretical results, supported by rigorous proofs, form the foundation of the proposed quantum stochasticity augmentation framework. They demonstrate how quantum noise, while inherent in NISQ devices, can be effectively managed to maintain robust and consistent representations. By leveraging these contributions, the proposed framework achieves enhanced generalization and accuracy for WiFi-based indoor localization under noisy conditions.
IV-A Theorem and Proof for Impact of Quantum Noise
Theorem IV.1.
(Similarity Consistency Under Quantum Noise): Cosine similarity between strong and weak augmented views remains consistent even when quantum noise affects the strong augmented view, thus it satisfies the condition shown in Eq. 57:
$\text{sim}(C_{st}^{i},C_{wt}^{i})=\text{sim}((C_{st}^{i}+\mathcal{N}_{st}),C_{wt}^{i}).$ (57)
Proof.
Let $C_{st}^{i}$ represent the strong augmented view of a quantum state, where the classical data is transformed into quantum states via quantum operations. Similarly, let $C_{wt}^{i}$ represent the weak augmented view, which remains unaffected by noise, then quantum noise $\mathcal{N}_{st}$, representing bit-flip, phase-flip, or depolarizing noise, is applied to the strong augmented view. The corrupted strong augmented view is defined and shown in Eq. 58.
$C_{st}^{i}=C_{st}^{i}+\mathcal{N}_{st}$ (58)
where $\mathcal{N}_{st}$ is the noise operator applied to the quantum state.
The cosine similarity between the strong and weak augmented views is used as the measure of similarity:
$\text{sim}(A,B)=\frac{A\cdot B}{\|A\|\|B\|}$ (59)
where $A\cdot B$ denotes the dot product, and $\|A\|,\|B\|$ are the Euclidean norms of the vectors $A$ and $B$.
For augmented views $C_{st}^{i}$ and $C_{wt}^{i}$, the similarity becomes:
$\text{sim}(C_{st}^{i},C_{wt}^{i})=\frac{C_{st}^{i}\cdot C_{wt}^{i}}{\|C_{st}^{i}\|\|C_{wt}^{i}\|}$ (60)
When quantum noise $\mathcal{N}_{st}$ is applied, the similarity is updated as:
$\text{sim}((C_{st}^{i}+\mathcal{N}_{st}),C_{wt}^{i})=\frac{(C_{st}^{i}+\mathcal{N}_{st})\cdot C_{wt}^{i}}{\|C_{st}^{i}+\mathcal{N}_{st}\|\|C_{wt}^{i}\|}$ (61)
Expanding the numerator:
$(C_{st}^{i}+\mathcal{N}_{st})\cdot C_{wt}^{i}=C_{st}^{i}\cdot C_{wt}^{i}+\mathcal{N}_{st}\cdot C_{wt}^{i}$ (62)
Given that quantum noise operators such as bit-flip, phase-flip, or depolarizing noise act orthogonally to the original quantum state, the term $\mathcal{N}_{st}\cdot C_{wt}^{i}$ approaches zero from IV.2.
$\mathbb{E}[\mathcal{N}_{st}\cdot C_{wt}^{i}]=0$ (63)
Thus, the similarity simplifies to:
$\text{sim}((C_{st}^{i}+\mathcal{N}_{st}),C_{wt}^{i})\approx\text{sim}(C_{st}^{i},C_{wt}^{i})$ (64)
Therefore, theroem is proved and can be written and shown in Eq. 65.
$\text{sim}(C_{st}^{i},C_{wt}^{i})=\text{sim}((C_{st}^{i}+\mathcal{N}_{st}),C_{wt}^{i})$ (65)
Therefore, the similarity remains balanced and consistent, even in the presence of quantum noise. This shows the resilience of our model to quantum noise (bit-flip, dephasing, and measurement noise), which overcomes issues related to noise in noisy environment. ∎
Lemma IV.2.
(Unbiased Quantum Noise Contribution): Quantum noise operators, being stochastic and uncorrelated with weak augmented views, introduce no systematic bias in the similarity computation.
$\mathbb{E}[\mathcal{N}_{st}\cdot C_{wt}^{i}]=0.$ (66)
Proof.
Let the weak augmented view $C_{wt}^{i}$ is a classical vector and remains unaffected by quantum noise. The strong augmented view $C_{st}^{i}$ is corrupted by a quantum noise operator $\mathcal{N}_{st}$, which represents the perturbations introduced by quantum processes such as bit-flip, phase-flip, or depolarizing noise. Then the first term, $C_{st}^{i}\cdot C_{wt}^{i}$, represents the expected similarity between the uncorrupted strong augmented view and the weak augmented view. The second term, $\mathcal{N}_{st}\cdot C_{wt}^{i}$, quantifies the effect of quantum noise on the similarity computation. To ensure that the similarity computation remains unbiased, we analyze the expectation of the second term:
$\mathbb{E}[\mathcal{N}_{st}\cdot C_{wt}^{i}]$ (67)
Assume that the quantum noise $\mathcal{N}_{st}$ is uncorrelated with the weak augmented view $C_{wt}^{i}$. This means that:
$\mathbb{E}[\mathcal{N}_{st}\cdot C_{wt}^{i}]=\mathbb{E}[\mathcal{N}_{st}]\cdot C_{wt}^{i}$ (68)
For typical quantum noise processes, such as bit-flip, phase-flip, or depolarizing noise, the expectation of the noise operator is zero due to its stochastic and unbiased nature:
$\mathbb{E}[\mathcal{N}_{st}]=0$ (69)
Substituting this result, we get:
$\mathbb{E}[\mathcal{N}_{st}\cdot C_{wt}^{i}]=0$ (70)
This shows that the quantum noise does not introduce systematic bias into the similarity computation, as the expected contribution of $\mathcal{N}_{st}\cdot C_{wt}^{i}$ is zero.
This completes the proof. ∎
V Empirical Validation and Discussion
This section provides a comprehensive evaluation of the proposed Quantum-Assisted Augmentation framework, highlighting its effectiveness in WiFi-based indoor localization tasks. By leveraging quantum-enhanced augmentations, the framework demonstrates significant improvements in robustness, generalization, and localization accuracy. The empirical results are validated on two benchmark datasets, UJIIndoorLoc [UJIIndoorLoc2014] and UTSIndoorLoc [UTSIndoorLoc2023], using a combination of contrastive learning and fine-tuning for regression and classification tasks, with performance evaluated against state-of-the-art methods for indoor localization.
V-A WiFi Datasets and Experimental Setup
The UJIIndoorLoc dataset contains 2,000 labeled instances collected across a university campus, featuring data from 52 WiFi access points. Each instance includes RSSI values, Floor_ID, Building_ID, and position identifiers. The dataset is evenly split into 1,000 training and 1,000 testing samples, with the primary objective being the prediction of building, floor, and device position based on WiFi signals. Preprocessing for this dataset involves mean imputation to handle missing RSSI values, min-max scaling to normalize signal strength within the range [0, 1], and one-hot encoding for categorical features such as Floor_ID and Building_ID. Instances with RSSI values below -90 dBm are discarded to reduce noise and enhance data quality. The UTSIndoorLoc dataset, collected within the University of Technology Sydney, comprises 5,000 instances of WiFi RSSI data from 38 WiFi access points across four floors. This dataset includes additional environmental variables, presenting a more complex scenario compared to UJIIndoorLoc. Similar preprocessing steps are applied: mean imputation addresses missing RSSI values, min-max normalization ensures uniform signal strength ranges, and one-hot encoding is used for Floor_ID and Building_ID. To mitigate the impact of unreliable measurements, instances with RSSI values below -90 dBm are removed.
Experimentation is conducted on a hardware setup comprising a PC equipped with an Intel i9 CPU operating at 3.20 GHz, 32GB of DDR5 random access memory, and an RTX 4090 GPU. This setup ensures efficient training and evaluation of the deep learning models, particularly for the high computational demands of the contrastive learning and quantum-enhanced augmentation processes. The model implementation is executed using TensorFlow and qiskit with GPU acceleration for optimized performance during training and inference.
Figure 4: Propose framework implementation result stepwise.
V-B Implementation Framework and Evaluation Results
This section outlines the implementation framework and evaluation process for the proposed method as shown in fig 4. The evaluation consists of data augmentation, training using contrastive learning, fine-tuning, and performance assessment. The input data undergoes quantum-assisted augmentation, generating two views—original and augmented (View 1 and View 2)—to simulate real-world variability and enhance the model’s generalization ability. These views are processed through the contrastive learning network, which optimizes the representations by clustering similar views and separating dissimilar ones in the latent space. The learned latent representations are visualized in two dimensions during the training and fine-tuning phases. Initially, the latent space after contrastive learning shows scattered points, indicating the model’s ability to map complex input data into a structured lower-dimensional space. Fine-tuning further refines this latent space, resulting in more distinct clusters, essential for accurate multi-floor and building localization. Performance is evaluated using a cumulative distribution function (CDF) curve of the distance error on the test set. The CDF illustrates the probability of achieving specific distance errors, with a steep rise indicating low errors with high probability. The proposed framework demonstrates superior localization accuracy, as evidenced by a sharp increase in the CDF curve, outperforming traditional methods. These results validate the robustness and effectiveness of the model in accurately predicting locations in diverse indoor environments.
Method UJIIndoorLoc Dataset UTSIndoorLoc Dataset
Building Accuracy Floor AccuracyMean Location Error (m) Floor Accuracy % Mean Location Error (m)
Random Forest 99.95 79.92 16.73 89.17 8.71
Decision Tree 98.91 78.03 22.89 77.31 12.69
KNN[Kim2018] 99.82 92.16 12.66 98.91 8.58
DNN[Kim2018] 99.10 79.12 17.53 97.16 9.26
CNNLoc[Song2019] 99.27 96.03 11.78 (34.85) 94.57 7.60 (9.77)
HADNNh[Cha2022] 99.82 93.88 11.27 87.11 10.62
HADNN[Cha2022] 100 93.15 11.59 92.01 10.36
CFNN[Cha2022] 100 92.34 11.74 89.17 10.32
GNN (baseline)[BaselineGNNWiFi_Wang2024] 100 94.15 9.61 97.91 7.48
Proposed 99.37 $\pm$ 0.63 96.89 $\pm$ 0.82 6.29 $\pm$ 0.57 98.12 $\pm$ 0.783 4.635 $\pm$ 0.825
TABLE I: Comparison of Proposed Approach with State-of-the-Art Works on Two Public Datasets
AP Selection Method Position Method UJIIndoorLoc Dataset UTSIndoorLoc Dataset
Building Accuracy (%) Floor Accuracy (%) Mean Location Error (m) Floor Accuracy (%) Mean Location Error (m)
AP Selection Network Our Approach 99.37 $\pm$ 0.63 96.89 $\pm$ 0.82 6.29 $\pm$ 0.57 98.12 $\pm$ 0.78 4.64 $\pm$ 0.83
GNN[BaselineGNNWiFi_Wang2024] 100 92.82 9.61 97.19 7.87
CFNN[Cha2022] 99.82 91.01 10.97 87.62 9.72
HADNN[Cha2022] 100 91.54 11.05 91.75 9.17
CNNLoc[Song2019] 99.87 92.80 35.67 94.08 10.29
AP Threshold Selection Our Approach 99.21 $\pm$ 0.42 96.43 $\pm$ 0.59 7.68 $\pm$ 0.64 97.58 $\pm$ 0.49 5.85 $\pm$ 0.67
GNN[BaselineGNNWiFi_Wang2024] 100 91.01 11.58 96.91 9.47
CFNN[Cha2022] 100 92.57 11.97 87.11 10.64
HADNN[Cha2022] 99.46 90.01 12.01 90.72 11.12
CNNLoc[Song2019] 99.82 91.56 36.68 91.78 11.50
AP Random Selection Our Approach 99.12 $\pm$ 0.59 96.26 $\pm$ 0.63 8.42 $\pm$ 0.87 97.08 $\pm$ 0.59 6.74 $\pm$ 0.26
GNN[BaselineGNNWiFi_Wang2024] 99.82 91.01 14.07 95.13 11.23
CFNN[Cha2022] 99.45 90.36 13.35 83.76 10.93
HADNN[Cha2022] 99.54 89.10 14.36 82.98 10.45
CNNLoc[Song2019] 99.75 88.98 38.44 91.25 11.18
TABLE II: Comparison Performance of Propose Work with S.O.T.A on different AP Selection Method for 70% of All APs on UJIIndoorLoc and UTSIndoorLoc Datasets
V-C Comprehensive Performance Comparison with Related Works
This section evaluates the proposed framework using the UJIIndoorLoc [UJIIndoorLoc2014] and UTSIndoorLoc [UTSIndoorLoc2023] datasets. The focus is on improving access point (AP) selection for multi-floor and building environments, with performance assessed using Building Accuracy, Floor Accuracy, and Mean Location Error. Tables I and II compare the proposed method with state-of-the-art (SOTA) approaches across different AP selection strategies.
V-C1 Performance on UJIIndoorLoc Dataset
The proposed method achieves 99.37% Building Accuracy, 96.89% Floor Accuracy, and a Mean Location Error of 6.29 meters. In comparison, the baseline GNN-based method[BaselineGNNWiFi_Wang2024] achieves perfect Building Accuracy (100%) but lags in Floor Accuracy (92.82%) and Mean Location Error (9.61 meters). CNNLoc[Song2019] shows significantly lower performance with a Mean Location Error of 35.67 meters, highlighting the limitations of traditional models in multi-floor environments.
V-C2 Performance on UTSIndoorLoc Dataset
For this dataset, the proposed method continues to excel with 98.12% Floor Accuracy and a Mean Location Error of 4.64 meters. The baseline GNN-based method[BaselineGNNWiFi_Wang2024] achieves 100% Building Accuracy but performs worse in Floor Accuracy (97.19%) and Mean Location Error (7.87 meters). These results underline the robustness of the proposed approach in diverse indoor environments.
V-C3 AP Selection Comparison
Table II compares the proposed method under three AP selection strategies: AP Selection Network, AP Threshold Selection, and AP Random Selection. Across all methods, the proposed approach consistently outperforms SOTA techniques. For example, with AP Threshold Selection on the UJIIndoorLoc dataset, it achieves 99.21% Building Accuracy, 96.43% Floor Accuracy, and a Mean Location Error of 7.68 meters, outperforming baseline GNN-based method[BaselineGNNWiFi_Wang2024], which exhibits a higher Mean Location Error (11.58 meters). Similarly, for AP Random Selection, the proposed method achieves 96.26% Floor Accuracy and 8.42 meters Mean Location Error, outperforming baseline GNN-based method[BaselineGNNWiFi_Wang2024] , which shows a Mean Location Error of 14.07 meters and others.
V-C4 Discussion:
Across both datasets and all AP selection strategies, the proposed framework demonstrates superior Building and Floor Accuracy while maintaining consistently low Mean Location Errors. The quantum-assisted data augmentation enhances robustness against noise and variability, providing reliable localization even in challenging scenarios. These results establish the proposed method as a robust solution for WiFi-based indoor localization, with strong potential for deployment in real-world multi-floor and multi-building environments.
Figure 5: CDA analysis and comparison of our approach with related works on a) UJIIndoorLoc dataset and b) UTSIndoorLoc dataset.
V-D CDF Analysis and Comparison
Figure 5 illustrates the cumulative distribution function (CDF) of the distance error for the UJIIndoorLoc and UTSIndoorLoc datasets. The analysis focuses on evaluating the effectiveness of the Proposed Work in minimizing localization errors compared to state-of-the-art models.
V-D1 UJIIndoorLoc Dataset
The CDF plot in Figure 5-a shows the distance error distribution for the UJIIndoorLoc dataset. The Proposed Work achieves the steepest curve closest to the y-axis, indicating consistently lower distance errors. The baseline GNN-based method[BaselineGNNWiFi_Wang2024] follows with competitive performance, while CFNN and HADCNN[Cha2022] exhibit higher errors, particularly at larger distance ranges. CNNLoc [Song2019] shows the worst performance, with a slower rise in the CDF curve, reflecting significant errors across the dataset.
V-D2 UTSIndoorLoc Dataset
Figure 5-b depicts the distance error distribution for the UTSIndoorLoc dataset. Similar to the UJIIndoorLoc dataset, the Proposed Work demonstrates the lowest distance error, outperforming all other models. The gap between the Proposed Work and the baseline GNN-based method[BaselineGNNWiFi_Wang2024] is narrower compared to the UJIIndoorLoc dataset but still favors the Proposed Work. CFNN and HADCNN[Cha2022] again show higher errors, with CNNLoc [Song2019] exhibiting the largest errors even at lower distance ranges.
V-D3 Discussion
Across both datasets, the Proposed Work consistently minimizes distance errors, as evidenced by its superior CDF curves. While the baseline GNN-based method[BaselineGNNWiFi_Wang2024] shows competitive results, other methods, including CFNN and HADCNN[Cha2022], and CNNLoc [Song2019], trail significantly in accuracy. These findings underscore the robustness and effectiveness of the Proposed Work in achieving high-precision indoor localization, particularly in challenging multi-floor environments.
Figure 6: Impact of Labeled Data w.t.t AP Selection, AP Threshold, and AP Random for UJIIndoorLoc and UTSIndoorLoc Datasets.
V-E Impact of Labeled Data
Figure 6 evaluates the influence of labeled data percentage on Mean Location Error and classification accuracy across different AP selection methods (AP Selection, AP Threshold, and AP Random) for the UJIIndoorLoc and UTSIndoorLoc datasets.
V-E1 Classification Accuracy
Figure 6-a shows the impact of labeled data percentage on classification accuracy for floor and building tasks. A positive correlation is observed between labeled data percentage and accuracy, with all methods benefiting from more labeled data. The AP Selection method consistently achieves the highest accuracy across both datasets, reflecting its ability to maximize the utility of labeled data for classification. The AP Threshold method demonstrates steady improvements, particularly for building and floor classification tasks. Meanwhile, the AP Random method, while showing improvements with more labeled data, lags behind the other two methods in terms of accuracy growth.
V-E2 Mean Location Error
Figure 6-b illustrates the relationship between labeled data percentage and Mean Location Error (in meters). As the percentage of labeled data increases, the error consistently decreases across all methods, emphasizing the importance of labeled data for fine-tuning. The AP Selection method achieves the lowest error on both datasets, particularly at higher labeled data percentages, showcasing its efficiency in reducing localization errors. The AP Threshold method also performs well, showing a steady reduction in error, albeit at a slower rate. Conversely, the AP Random method exhibits higher error rates, indicating that random AP selection is less effective in leveraging labeled data.
V-E3 Dataset Comparison
The UJIIndoorLoc dataset slightly outperforms UTSIndoorLoc in both building and floor classification accuracy, indicating that it may have more structured or simpler data conducive to better model performance. For both datasets, the AP Selection method remains the most effective for minimizing Mean Location Error and maximizing classification accuracy.
V-E4 Discussion
These findings highlight the critical role of labeled data in enhancing model performance. The AP Selection method emerges as the most effective strategy, consistently reducing location error and improving classification accuracy. Fine-tuning with labeled data significantly benefits the proposed framework, with clear improvements observed in both metrics as labeled data increases. However, the AP Random method’s slower rate of improvement underscores the value of targeted selection methods like AP Selection and AP Threshold for WiFi-based indoor localization tasks.
VI Conclusion
This paper introduced Quantum Stochastic Contrastive Learning (QSCL), a novel framework for WiFi-based indoor localization designed to address key challenges such as data variability, environmental noise, and limited labeled data. Grounded in rigorous theoretical foundations, QSCL incorporates four theorems and one lemma, formally proving its capabilities in probabilistic augmentation, diversity enhancement, relationship preservation, and resilience to quantum noise. By leveraging quantum computing for strong data augmentations with stochastic perturbations, and combining these with classical weak augmentations, QSCL enhances data diversity and facilitates robust feature learning. The framework integrates a Spatio-Temporal Adaptive Attention (STAA) encoder to effectively capture spatial and temporal dependencies in sequential data. Additionally, a bidirectional contrastive loss function ensures robust representations by capturing forward and reverse relationships between augmented views. Evaluations on the UJIIndoorLoc and UTSIndoorLoc datasets demonstrate the framework’s superior performance in localization accuracy, floor and building detection, and generalization, even under noisy conditions. QSCL achieves significant improvements while reducing reliance on labeled data and exhibiting resilience to quantum noise, including bit-flip, dephasing, and measurement noise. This work advances the state-of-the-art in WiFi-based indoor localization and highlights the potential of integrating quantum computing with machine learning to tackle complex challenges in real-world applications. Future work will focus on implementing QSCL on real quantum hardware, exploring scalability to diverse environments, and integrating multimodal data sources to further enhance localization accuracy and robustness.
[Uncaptioned image] Muhammad Bilal Akram Dastagir Muhammad Bilal Akram Dastagir is a PhD graduand in Computer Science at the School of Computing at the Korea Advanced Institute of Science and Technology (KAIST) in Daejeon, South Korea, His academic journey and professional expertise encompass developing novel algorithms and architectures that integrate quantum mechanics with artificial intelligence, focusing on real-world applications such as Indoor-Outdoor Detection, Floor Change Detection, and Deep Learning for WiFi-based Indoor Localization Solutions. and IoT security. Before this, he worked as a manager (Cryptography Engineer) for many years at the National Space Agency, SUPARCO, and was awarded the Service Distinguished Award by Chairman SUPARCO, for his contribution to the design and development of the security protocol for Telecomands and Telemetry. Prior, he completed his Master’s degree in Advanced Engineering at the University of Leicester in the United Kingdom.
[Uncaptioned image] Dongsoo Han (Member, IEEE) received the Ph.D. degree in information science from Kyoto University. He is currently a Professor of Computer Science at the Korea Advanced Institute of Science and Technology (KAIST), where he is also the Director of the Indoor Positioning Research Center. His research interests include indoor positioning, pervasive computing, and location-based mobile applications.
[Uncaptioned image] Saif Al-Kuwari received the Bachelor of Engineering in computers and networks from the University of Essex, Colchester, U.K., in 2006, the first Ph.D. degree in computer science from the University of Bath and Royal Holloway, Egham, Surrey, in 2011, and the second Ph.D. degree in computer science from the University of London, London, U.K., in 2011. He is currently an Assistant Professor with the College of Science and Engineering, Hamad Bin Khalifa University, Qatar Foundation, Doha, Qatar. His research interests include quantum computing, cryptography, computational forensics, and their connections with machine learning. Dr. Al-Kuwari is an IET and BCS Fellow, and an ACM Senior Member.
[Uncaptioned image] Ahmed Farouk is currently an Assistant Professor with South Valley University, Hurghada, Egypt. Before that, he was a Postdoctoral Research Fellow with Wilfrid Laurier University and Ryerson University, Canada. He has been awarded the Lindau Nobel Laureate Alumni, CDL University of Toronto Alumni, and Outstanding IEEE Computer Chapter for K-W Region 2020. He is exceptionally well known for his seminal contributions to theories of quantum information, machine learning, and cryptography. He has published over 80 articles in reputed and high-impact publications like the IEEE Internet of Things Journal, IEEE Transactions on Intelligent Transportation Systems, IEEE Wireless Communications, and IEEE Transactions on Industrial Informatics (IEEE TII). His volunteer work is apparent since he was appointed as the Chair of the IEEE Computer Society Chapter for the Waterloo-Kitchener area and has joined the editorial boards of many reputed journals. Recently, he was appointed as an Officer (Secretary) of the IEEE Technical Committee on Quantum in Consumer Technology.
[…]